Ali walks 440m east, 250m south 80m east. Find the distance and displacement
1
Expert's answer
2020-03-16T13:16:42-0400

Distance:


"d=440+250+80=770\ m"

Displacement:


"s=\sqrt{(440+80)^2+250^2}=577\ m"

Direction:


"\theta=\arctan{\frac{250}{520}}=25.7\degree\text{S of E}"